A loan has an initial balance of $100,000 and has interest charged at 8% compounded annually. You are going to pay the loan back with even payments in years 1, 2, 3, 4, 6, 7, 8, and 9. Find the payment amount.

$17965

$15965

$11965

$13965
Answer #1

Let annual loan payment amount be A, then as per given condition

A *(P/A,8%,4) + A *(P/A,8%,4)*(P/F,8%,5) = 100000

A *3.312127 + A * 3.312127*0.680583 = 100000

A *5.5663043 = 100000

A = 100000 / 5.5663043

A = 17965.24 ~ 17965

First option is correct answer